what is the answer for log(2x+1) = 1 + log (x-2) ?
Found 2 solutions by stanbon, josgarithmetic:
Answer by stanbon(75887)

what is the answer for
log(2x+1) = 1 + log (x-2)
----
log(2x+1) - log(x-2) = 1
-----
log[(2x+1)/(x-2)] = 1
-----
(2x+1)/(x-2) = 10^1
----
Cross-multiply
2x+1 = 10(x-2)
2x + 1 = 10x -20
8x = 21
x = 21/8
=================
Cheers,
Stan H.
================

Answer by josgarithmetic(39621)

Steps:
log(2x+1)-log(x-2)=1
log((2x+1)/(x-2))=1 [property of logarithms]

You need to know which base to continue. Is the base 10, or 2, or e? If base is 10, then (2x+1)/(x-2)=10,
2x+1=10x-20,
1=8x-20
-19=8x,
x=-19/8
Is base 2?
(2x+1)=2(x-2)
2x+1=2x-4
False
Is the base e?